OpenStudy (anonymous):

Jackie's report is 5 pages longer than half the number in Kane's Report.There are 14 pages in Kane's report

OpenStudy (anonymous):

J = 1/2k + 5 K=14 J= 1/4(14)+5 =7+5 J=12

OpenStudy (anonymous):

Im meant 1/2 instead of 1/4 btw

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

hey how old are u

OpenStudy (anonymous):

The Length of a rectangle is 4 times it's width.The Rectangle's width is 8 m